Using the limit definition compute the instantaneous rate of change of the function at x=a f(x)=9x^2, a=8
answer is 144, I keep getting 153. Please help!

Answer by CharlesG2(834)

Using the limit definition compute the instantaneous rate of change of the function at x=a f(x)=9x^2, a=8
answer is 144, I keep getting 153. Please help!
(not sure if you put this in right category, but nevermind)
Instantaneous Rate of Change --> The rate of change at a particular moment. Same as the value of the derivative at a particular point.


f(x) = 9x^2, x = a, where a = 8
f'(x) = 2 * 9 * x = 18x (took derivative)
f'(a) = 18a
f'(8) = 18 * 8 = 144
have no idea how you got 153